# Learning to Optimize for Mixed-Integer Non-linear Programming

### v3

@@ Methods.
Table 1 provides an overview of all the methods used in the following experiments. A 60-second time limit is enforced for all methods and problems. The experiments evaluate our learning-based methods, Rounding Classification (RC) and Learnable Threshold (LT), against traditional exact optimization (EX), which can compute optimal solutions but is often computationally expensive, and heuristic-based approaches such as Rounding after Relaxation (RR) and root node solutions (N1), which offer faster results without quality guarantees. Note that baselines EX and N1 include a wide range of heuristics that are embedded in the MINLP solver of choice (Gurobi or SCIP) and that are executed in conjunction with the tree search procedure; we are also implicitly comparing to these heuristics, not just to the exact search. As such, the competing methods cover a broad spectrum of optimization strategies, from exact solvers to fast heuristics, allowing for a comprehensive evaluation of solution quality and computational efficiency. In addition, two ablation baselines are included: Rounding after Learning (RL), which applies direct rounding to the relaxed solution generated by the learning-based method without any corrective adjustments, and Rounding with STE (RS), which uses the Straight-Through Estimator to handle non-differentiable rounding but without a learning-based correction layer.
@@
Rounding after Learning* RL* Learning-based method that learns from the relaxed problem, then rounds the inferred solution to the nearest integer.
Rounding with STE* RS* Learning-based method where the rounding process is differentiable with STE but does not involve learning.
@@
Root Node Solution N1 A feasible solution obtained at the root node, which uses heuristics after continuous relaxation with cutting planes.
@@
We tested the methods on a variety of optimization problems, including integer convex quadratic problems, simple integer non-convex problems, and high-dimensional mixed-integer Rosenbrock problems. These problem classes were selected to cover both convex and non-convex scenarios and to evaluate the scalability of the methods in higher-dimensional settings. Each method was assessed in terms of objective value, constraint violation, and solving time, providing a comprehensive view of their performance across different types of problems. In addition, we performed experiments on integer linear programs in Appendix E; these experiments simply serve to show that our methods could be applied to the integer linear case, though the use of MILP solvers may be preferable.
@@
The solution mapping $\pi_{\Theta_{1}}$ used across all learning-based methods (RC, LT, and RL) and the rounding correction network $\varphi_{\Theta_{2}}$ for RC and LT are based on fully connected layers with ReLU activations. Further details regarding the network hyperparameters can be found in Appendix B. For all problems, the training samples 8,000 instances from the distribution, and the test set includes 100 instances. An additional set of 1,000 instances was used for validation to fine-tune the models and select hyperparameters.
@@
All experiments were conducted on a system with an AMD Ryzen 7 5800HS CPU, 15GB of DDR4 RAM, and an NVIDIA GeForce GTX 1650 GPU. The software environment was configured with Python 3.10.13, PyTorch 2.0.1+cu117 (Paszke et al., 2019) for deep learning models, and NeuroMANCER 1.5.1 (Drgona et al., 2023) for modeling parametric constrained optimization problems.
Gurobi 11.0.3 (Gurobi Optimization, LLC, 2021) is used as the exact method for convex quadratic problems; beyond quadratic polynomials, Gurobi needs to approximate non-linearities using piecewise-linear functions. For those more general mixed-integer non-convex problems, we use SCIP 8.1.0 (Bestuzheva et al., 2021) with Ipopt 3.14.13 (Wächter & Biegler, 2006) as the continuous non-linear solver. Note that Gurobi and SCIP are considered to be among the state-of-the-art solvers for MINLP, as noted by Lundell & Kronqvist (2022) who performed a comprehensive benchmarking of more than ten MINLP solvers: “It is clear, however, that the global solvers Antigone, BARON, Couenne and SCIP are the most efficient at finding the correct primal solution when regarding the total time limit. […] Gurobi also is very efficient when considering that it only supports a little over half of the total number of problems!”
@@ 6.1 Convex Quadratic Problem
Table 2: Result for a Convex Quadratic Problem. Each problem size is evaluated on a test set of 100 instances. “Obj Mean” and “Obj Median” represent the mean and median objective values for this minimization problem, with smaller values being better. “Infeasible” denotes the fraction of infeasible solutions, and “Time (Sec)” is the average solving/inference time per instance. The “—” symbol indicates that no solution is found for any instance within 60 sec.
Method Metric 5×5 10×10 20×20 50×50 100×100 200×200 500×500
RC Obj Mean $0.608$ $-1.602$ $-4.148$ $-13.109$ $-13.589$ $-31.206$ $-73.570$
Obj Median $0.413$ $-1.600$ $-4.194$ $-13.131$ $-13.587$ $-31.247$ $-73.761$
% Infeasible $3\%$ $2\%$ $4\%$ $1\%$ $4\%$ $5\%$ $8\%$
Time (Sec) $0.0026$ $0.0031$ $0.0038$ $0.0031$ $0.0039$ $0.0044$ $0.0074$
LT Obj Mean $0.649$ $-1.818$ $-4.329$ $-12.905$ $-13.745$ $-31.261$ $-73.457$
Obj Median $0.470$ $-2.078$ $-4.341$ $-12.946$ $-13.747$ $-31.320$ $-73.414$
% Infeasible $5\%$ $7\%$ $1\%$ $2\%$ $7\%$ $1\%$ $8\%$
Time (Sec) $0.0032$ $0.0030$ $0.0038$ $0.0031$ $0.0045$ $0.0045$ $0.0098$
EX Obj Mean $0.294$ $-2.779$ $-5.120$ $-15.885$
Obj Val Median $0.129$ $-2.991$ $-5.130$ $-15.940$
% Infeasible $0\%$ $0\%$ $0\%$ $0\%$
Time (Sec) $0.1451$ $0.1996$ $5.4154$ $61.2890$
RR Obj Mean $0.211$ $-2.858$ $-5.179$ $-16.173$ $-21.922$ $-46.727$ $-106.526$
Obj Median $0.058$ $-3.033$ $-5.217$ $-16.205$ $-21.892$ $-46.755$ $-106.536$
% Infeasible $97\%$ $100\%$ $100\%$ $100\%$ $100\%$ $100\%$ $100\%$
Time (Sec) $0.1316$ $0.1398$ $0.1307$ $0.3320$ $0.1697$ $0.3288$ $1.4590$
N1 Obj Mean $0.759$ $151.2e$ 15 $104.4e$ 10 $171.7e$ 17 $181.7e$ 18
Obj Median $0.455$ $-2.033$ $9.581$ $172.4e$ 17 $181.5e$ 18
@@
Time (Sec) $0.1301$ $0.1322$ $0.1316$ $0.2810$ $43.9388$
RL Obj Mean $0.493$ $-2.481$ $-4.660$ $-14.155$ $-17.195$ $-37.761$ $-90.205$
Obj Median $0.301$ $-2.756$ $-4.685$ $-14.163$ $-17.149$ $-37.761$ $-90.205$
% Infeasible $58\%$ $93\%$ $52\%$ $44\%$ $82\%$ $95\%$ $100\%$
Time (Sec) $0.0007$ $0.0007$ $0.0007$ $0.0011$ $0.0008$ $0.0009$ $0.0008$
RS Obj Mean $0.941$ $-1.097$ $-3.945$ $-11.982$ $-11.003$ $-24.598$ $-57.312$
Obj Median $0.777$ $-1.409$ $-3.956$ $-12.002$ $-11.205$ $-24.566$ $-57.312$
% Infeasible $0\%$ $0\%$ $0\%$ $0\%$ $0\%$ $0\%$ $0\%$
Time (Sec) $0.0014$ $0.0018$ $0.0020$ $0.0015$ $0.0020$ $0.0043$ $0.0038$
The convex quadratic problems used in the experiments are based on Donti et al. (2021), which also experimented with learning under hard (but not integer) constraints. Several modifications were introduced to ensure compatibility with our discrete version. First, all the decision variables all restricted to be integers. Second, the coefficients were rescaled. Furthermore, the original inequality constraints were removed, and the original equality constraints were transformed into inequalities:
@@
where the coefficients $\bm{Q}\in\mathbb{R}^{n\times n}$, $\bm{p}\in\mathbb{R}^{n}$, and $\bm{A}\in\mathbb{R}^{m\times n}$ were fixed and generated using random values, while $\bm{b}\in\mathbb{R}^{m}$ were treated as parametric coefficients (input features), varying across instances. Further details on the generation of $\bm{Q},\bm{p},\bm{A},$ and $\bm{b}$ can be found in Section D.1.
We experimented with quadratic problems of different sizes, from $5$ decision variables and $5$ constraints ($5\times 5$) up to $500\times 500$. The results in Table 2 summarize the performance of all methods across different problem sizes. The RC and LT methods exhibit robust performance across the board, achieving objective values second only to EX while consistently maintaining low percentages of infeasible solutions and fast solution times across all problem sizes. These methods achieve several orders of magnitude speed-ups, scaling effectively even for large instances up to $500\times 500$. Without learning-based correction, the objective values of RS are generally worse than those of RC and LT, but there are no infeasible solutions. The exact solver EX, while performing well on smaller problem sizes, fails to produce any solutions for instances of size $100\times 100$ and larger within the 60-second time limit, highlighting its limitations when handling more complex problems. N1, on the other hand, can find feasible solutions within a short time frame for smaller cases but suffers from severe numerical instability as the problem size increases. When scaled to $200\times 200$, N1 also fails to produce a solution. Both the RR and RL methods, which rely on rounding relaxations of either the solver or the prediction model, encounter significant feasibility challenges. Overall, this analysis underscores that learning-based methods like RC and LT offer considerable advantages in both solution quality and computational speed, especially for large-scale problems, compared to exact solvers or other heuristics.
@@
In addition to evaluating solution quality, feasibility, and solving/inference times, we also measured the offline training times for our two approaches on different problem sizes. These results, along with training times for other problem types, are presented in Appendix F, where it is evident that the training times for the learning-based methods scale well with problem size.
@@ 6.2 Simple Non-convex Problem
Table 3: Results for a Simple Non-convex Problem. See the caption of Table 2 for details. “Unsolved” denotes the percentage of instances that EX or N1 could not solve within 60 seconds and is omitted for other methods, which either solve all or none.
Method Metric 5×5 10×10 20×20 50×50 100×100 200×200 500×500
RC Obj Mean $0.202$ $0.988$ $0.397$ $0.678$ $0.974$ $1.688$ $-0.716$
Obj Median $0.163$ $0.676$ $0.384$ $0.649$ $0.928$ $1.688$ $-0.717$
% Infeasible $0\%$ $2\%$ $1\%$ $2\%$ $1\%$ $0\%$ $4\%$
Time (Sec) $0.0024$ $0.0028$ $0.0037$ $0.0025$ $0.0036$ $0.0032$ $0.0061$
LT Obj Mean $0.211$ $0.605$ $-0.000$ $0.424$ $1.038$ $0.514$ $-1.646$
Obj Median $0.143$ $0.349$ $-0.018$ $0.410$ $1.014$ $0.464$ $-1.656$
% Infeasible $0\%$ $3\%$ $0\%$ $0\%$ $1\%$ $1\%$ $0\%$
Time (Sec) $0.0030$ $0.0036$ $0.0035$ $0.0033$ $0.0032$ $0.0033$ $0.0068$
EX Obj Mean $-0.027$ $-0.142$ $1.936$ $18.874$
Obj Median $-0.058$ $-0.317$ $0.288$ $11.729$
% Infeasible $0\%$ $2\%$ $2\%$ $11\%$
% Unsolved $0\%$ $3\%$ $73\%$ $77\%$
Time (Sec) $0.2046$ $4.0529$ $18.7699$ $19.5122$
RR Obj Mean $-0.047$ $-0.149$ $27.721$ $-0.782$
Obj Median $-0.089$ $-0.319$ $-0.3481$ $-0.711$
% Infeasible $63\%$ $87\%$ $98\%$ $100\%$
Time (Sec) $0.7142$ $1.5372$ $50.6019$ $54.4754$
N1 Obj Mean $1.272$ $4.651$ $29.918$ $289.992$
Obj Median $0.286$ $0.371$ $0.948$ $17.727$
% Infeasible $0\%$ $2\%$ $2\%$ $11\%$
% Unsolved $0\%$ $2\%$ $1\%$ $57\%$
Time (Sec) $0.0926$ $1.5369$ $1.8148$ $11.4497$
RL Obj Mean $0.179$ $3.282$ $5.158$ $16.626$ $19.653$ $41.756$ $93.030$
Obj Median $0.157$ $3.079$ $5.123$ $16.785$ $19.591$ $41.756$ $93.030$
% Infeasible $58\%$ $93\%$ $52\%$ $44\%$ $78\%$ $95\%$ $100\%$
Time (Sec) $0.0008$ $0.0006$ $0.0009$ $0.0010$ $0.0010$ $0.0010$ $0.0009$
RS Obj Mean $0.237$ $3.819$ $4.694$ $15.280$ $21.124$ $40.595$ $104.886$
Obj Median $0.186$ $3.503$ $4.730$ $15.156$ $20.874$ $40.639$ $104.886$
% Infeasible $0\%$ $0\%$ $0\%$ $0\%$ $0\%$ $0\%$ $0\%$
Time (Sec) $0.0019$ $0.0027$ $0.0021$ $0.0026$ $0.0020$ $0.0022$ $0.0040$
@@
where the sine function is applied element-wise to the decision variables $\bm{x}$. This introduces non-convexity into the problem, making it more challenging compared to the convex case. The coefficients $\bm{Q}$, $\bm{p}$, $\bm{A}$, and $\bm{b}$ are generated in the same way as in the quadratic formulation (Section D.1). In addition, the scales of the problem and the experiment setting are also identical to those of the quadratic problems.

### v4

@@ Methods.
Table 1 provides an overview of all the methods used in the following experiments. A 60 or 1000-second time limit is enforced for all methods and problems. The experiments evaluate our learning-based methods, Rounding Classification (RC) and Learnable Threshold (LT), against traditional exact optimization (EX), which can compute optimal solutions but is often computationally expensive, and heuristic-based approaches such as Rounding after Relaxation (RR) and root node solutions (N1), which offer faster results without quality guarantees. Note that baselines EX and N1 include a wide range of heuristics that are embedded in the MINLP solver of choice (Gurobi or SCIP) and that are executed in conjunction with the tree search procedure; we are also implicitly comparing to these heuristics, not just to the exact search. As such, the competing methods cover a broad spectrum of optimization strategies, from exact solvers to fast heuristics, allowing for a comprehensive evaluation of solution quality and computational efficiency. In addition, we evaluate two ablation baselines, which isolate different aspects of our correction layers $\varphi_{\Theta_{2}}$ to highlight their impact on performance. Details of these ablation studies, including methodology and results, are provided in Appendix F.
@@
Root Node Solution N1 A feasible solution from the root node of the solver, which uses heuristics after continuous relaxation with cutting planes.
@@
We tested the methods on a variety of optimization problems, including integer convex quadratic problems, simple integer non-convex problems, and high-dimensional mixed-integer Rosenbrock problems. These problem classes were selected to cover both convex and non-convex scenarios and to evaluate the scalability of the methods in higher-dimensional settings. Each method was assessed in terms of objective value, constraint violation, and solving time, providing a comprehensive view of their performance across different types of problems. In addition, we evaluated our methods on integer linear programs (MILPs), in which the dataset from the MIP Workshop 2023 Computational Competition Bolusani et al. (2023). These experiments primarily serve to demonstrate that our methods can also handle integer linear cases, though the use of MILP solvers may be preferable. Further details are provided in Appendix H.
@@
The solution mapping $\pi_{\Theta_{1}}$ used across all learning-based methods (RC, LT, and RL) and the rounding correction network $\varphi_{\Theta_{2}}$ for RC and LT are based on fully connected layers with ReLU activations. Further details regarding the network hyperparameters can be found in Appendix D. For all problems, the training samples 8,000 instances from the distribution, and the test set includes 100 instances. An additional set of 1,000 instances was used for validation to fine-tune the models and select hyperparameters.
@@
All experiments were conducted on a system with 2 Intel Silver 4216 Cascade Lake @ 2.1GHz CPUs, 64GB RAM, and 4 NVIDIA V100 Volta GPUs. The software environment was configured with Python 3.10.13, PyTorch 2.5.0+cu122 (Paszke et al., 2019) for deep learning models, and NeuroMANCER 1.5.2 (Drgona et al., 2023) for modeling parametric constrained optimization problems. Gurobi 11.0.1 (Gurobi Optimization, LLC, 2021) is used as the exact method for convex quadratic problems; beyond quadratic polynomials, Gurobi needs to approximate non-linearities using piecewise-linear functions. For those more general mixed-integer non-convex problems, we use SCIP 9.0.0 (Bestuzheva et al., 2021) with Ipopt 3.14.14 (Wächter & Biegler, 2006) as the continuous non-linear solver. Note that Gurobi and SCIP are considered to be among the state-of-the-art solvers for MINLP, as noted by Lundell & Kronqvist (2022) who performed a comprehensive benchmarking of more than ten MINLP solvers: “It is clear, however, that the global solvers Antigone, BARON, Couenne and SCIP are the most efficient at finding the correct primal solution when regarding the total time limit. […] Gurobi also is very efficient when considering that it only supports a little over half of the total number of problems!”
Overall results.
As shown in Figure 2, the exact solvers like Gurobi and SCIP gradually improve the objective value over time, but this often comes at a high computational cost. For more complex problems, they may even fail to find feasible solutions within reasonable time limits. In contrast, our methods, RC and LT, achieve high-quality, feasible solutions in mere milliseconds. Even when accounting for the 131.72 seconds required to train the neural network for the Rosenbrock problem, our approaches remain significantly more efficient. Once trained, these models generalize well to unseen instances, making them ideal for repeated problem-solving scenarios where the training cost is amortized. Additionally, RC and LT could provide high-quality initial solutions for exact solvers, reducing the search space and accelerating convergence, thus enhancing the performance of traditional optimization methods.
Figure 2: Illustration of objective value evolution for a $100\times 100$ Convex Quadratic instance and a $60\times 4$ Rosenbrock instance over 600 seconds.
@@ 6.2 Convex Quadratic Problem
Table 2: Result for a Convex Quadratic Problem. Each problem size is evaluated on a test set of 100 instances. “Obj Mean” and “Obj Median” represent the mean and median objective values for this minimization problem, with smaller values being better. “Infeasible” denotes the fraction of infeasible solutions, and “Time (Sec)” is the average solving/inference time per instance. The “—” symbol indicates that no solution is found for any instance within 1000 seconds.
Method Metric 5×5 10×10 20×20 50×50 100×100 200×200 500×500 1000×1000
RC Obj Mean $0.803$ $-1.476$ $-4.151$ $-12.422$ $-13.564$ $-30.799$ $-71.373$ $-120.378$
Obj Median $0.648$ $-1.719$ $-4.166$ $-12.416$ $-13.634$ $-30.789$ $-71.476$ $-120.492$
% Infeasible $0\%$ $3\%$ $4\%$ $0\%$ $4\%$ $1\%$ $4\%$ $3\%$
Time (Sec) $0.0019$ $0.0019$ $0.0019$ $0.0019$ $0.0021$ $0.0025$ $0.0026$ $0.0045$
LT Obj Mean $0.743$ $-1.615$ $-4.165$ $-12.355$ $-13.395$ $-29.728$ $-71.350$ $-114.434$
Obj Median $0.554$ $-1.928$ $-4.203$ $-12.352$ $-13.527$ $-29.808$ $-71.601$ $-114.522$
% Infeasible $1\%$ $1\%$ $0\%$ $4\%$ $3\%$ $0\%$ $3\%$ $1\%$
Time (Sec) $0.0019$ $0.0019$ $0.0019$ $0.0019$ $0.0023$ $0.0022$ $0.0026$ $0.0046$
EX Obj Mean $0.294$ $-2.779$ $-5.120$ $-15.928$ $-20.790$
Obj Median $0.129$ $-2.991$ $-5.130$ $-15.956$ $-20.778$
@@
Time (Sec) $0.496$ $0.664$ $8.728$ $1520.733$ $1237.534$
RR Obj Mean $0.211$ $-2.858$ $-5.179$ $-16.173$ $-21.922$ $-46.727$ $-106.526$ $-213.312$
Obj Median $0.058$ $-3.033$ $-5.217$ $-16.205$ $-21.892$ $-46.755$ $-106.536$ $-213.292$
% Infeasible $97\%$ $100\%$ $100\%$ $100\%$ $100\%$ $100\%$ $100\%$ $100\%$
Time (Sec) $0.411$ $0.412$ $0.417$ $0.440$ $0.583$ $0.846$ $2.639$ $8.874$
N1 Obj Mean $0.549$ $151.2e$ 15 $079.8e$ 07 $171.7e$ 17 $181.5e$ 18
Obj Median $0.369$ $-1.900$ $9.600$ $172.4e$ 17 $181.4e$ 18
% Infeasible $0\%$ $0\%$ $0\%$ $0\%$ $0\%$
Time (Sec) $0.420$ $0.422$ $0.415$ $0.498$ $104.204$
Since there is a lack of publicly available datasets for parametric MINLPs, the convex quadratic problems used in the experiments are adapted from Donti et al. (2021), which originally focused on learning under continuous constraints. We introduced integrality constraints on all decision variables to tailor these problems to our discrete setting. Additionally, we removed equality constraints to avoid the issue of generating infeasible instances. These modifications ensure compatibility with our framework while preserving the essential structure of the original problems:
@@
where the coefficients $\bm{Q}\in\mathbb{R}^{n\times n}$, $\bm{p}\in\mathbb{R}^{n}$, and $\bm{A}\in\mathbb{R}^{m\times n}$ were fixed, while $\bm{b}\in\mathbb{R}^{m}$ were treated as parametric coefficients (input features), varying across instances. Further details on the generation of $\bm{Q},\bm{p},\bm{A},$ and $\bm{b}$ can be found in Appendix E.
We experimented with quadratic problems of different sizes, from $5$ decision variables and $5$ constraints ($5\times 5$) up to ($1000\times 1000$). The results in Table 2 summarize the performance of all methods across different problem sizes. For a detailed analysis of constraint violation metrics, please refer to Appendix G. The RC and LT methods exhibit robust performance across the board, achieving objective values second only to EX while consistently maintaining low percentages of infeasible solutions and fast solution times across all problem sizes. These methods achieve several orders of magnitude speed-ups, scaling effectively even for large instances up to $1000\times 1000$. The exact solver EX, while performing well on smaller problem sizes, fails to produce any solutions for instances of size $200\times 200$ and larger within the 1000-second time limit, highlighting its limitations when handling more complex problems. N1, on the other hand, can find feasible solutions within a short time frame for smaller cases but suffers from severe numerical instability as the problem size increases. When scaled to $200\times 200$, N1 also fails to produce a solution. The RR method, which relies on rounding relaxations, encounters significant feasibility challenges. Overall, this analysis underscores that learning-based methods like RC and LT offer considerable advantages in both solution quality and computational speed, especially for large-scale problems, compared to exact solvers or other heuristics.
@@
In addition to evaluating solution quality, feasibility, and solving/inference times, we also measured the offline training times for our two approaches on different problem sizes. These results, along with training times for other problem types, are presented in Appendix I, where it is evident that the training times for the learning-based methods scale well with problem size.
@@ 6.3 Simple Non-convex Problem
where the sine function is applied element-wise to the decision variables $\bm{x}$. This introduces non-convexity into the problem, making it more challenging compared to the convex case. The constraint matrix $\bm{A}$ is parametrized by $\bm{d}$ (details in Appendix E), an additional variable that influences the matrix structure, further enriching the complexity. Alongside $\bm{d}$, the right-hand side vector $\bm{b}$ remains a dynamic parameter in the problem. The coefficients $\bm{Q}$, $\bm{p}$, $\bm{A}$, and $\bm{b}$ are generated in the same way as in the quadratic formulation. In addition, the scales of the problem and the experiment setting are also identical to those of the quadratic problems.
Table 3: Results for a Simple Non-convex Problem. See the caption of Table 2 for details. The “—” symbol indicates that no solution is found for any instance within 60 seconds.
Method Metric 5×5 10×10 20×20 50×50 100×100 200×200 500×500 1000×1000
RC Obj Mean $0.383$ $1.102$ $0.228$ $0.771$ $1.664$ $1.472$ $0.526$ $1.422$
Obj Median $0.240$ $0.838$ $0.217$ $0.752$ $1.594$ $1.436$ $0.526$ $0.809$
% Infeasible $1\%$ $3\%$ $0\%$ $2\%$ $0\%$ $1\%$ $4\%$ $3\%$
Time (Sec) $0.0019$ $0.0019$ $0.0019$ $0.0020$ $0.0022$ $0.0022$ $0.0029$ $0.0040$
LT Obj Mean $0.359$ $0.910$ $0.195$ $0.580$ $0.669$ $-0.356$ $-1.374$ $-3.744$
Obj Median $0.226$ $0.683$ $0.175$ $0.566$ $0.649$ $-0.373$ $-1.594$ $-3.716$
% Infeasible $0\%$ $2\%$ $1\%$ $2\%$ $4\%$ $0\%$ $2\%$ $1\%$
Time (Sec) $0.0019$ $0.0019$ $0.0019$ $0.0020$ $0.0021$ $0.0023$ $0.0029$ $0.0050$
EX Obj Mean $0.001$ $-0.182$ $-0.406$ $25.750$
Obj Median $-0.095$ $-0.310$ $-0.420$ $9.520$
% Infeasible $0\%$ $0\%$ $0\%$ $0\%$
Time (Sec) $0.118$ $2.509$ $60.153$ $60.125$
RR Obj Mean $-0.047$ $-0.168$ $-0.464$ $-1.039$ $-2.068$ $-3.990$ $-7.935$
Obj Median $-0.089$ $-0.325$ $-0.476$ $-1.215$ $-2.307$ $-4.327$ $-7.104$
% Infeasible $64\%$ $86\%$ $97\%$ $100\%$ $100\%$ $100\%$ $56\%$
Time (Sec) $0.216$ $0.422$ $1.013$ $1.198$ $4.654$ $52.242$ $62.006$
N1 Obj Mean $1.690$ $31.1e$ 3 $42.1e$ 4 $63.7e$ 6
Obj Median $0.183$ $0.557$ $2.222$ $45.847$
% Infeasible $0\%$ $0\%$ $0\%$ $0\%$
Time (Sec) $0.0434$ $0.1029$ $0.1516$ $8.936$